SUBFIG ONE

(a) Subfigure One.

SUBFIG TWO

(b) Subfigure Two.

Figure 1: Two side-by-side figures.

SUBFIG THREE

(a) Subfigure Three.

SUBFIG FOUR

(b) Subfigure Four.

SUBFIG FIVE

(c) Subfigure Five.

SUBFIG SIX

(d) Subfigure Six.

Figure 2: Four figures with specified suppression of extra padding.

Figures wrapped to show any extra spaces introduced in processing the subfigures.

SUBFIG SEVEN

(a) Subfigure Seven.

SUBFIG EIGHT

(b) Subfigure Eight.

SUBFIG NINE

(c) Subfigure Nine.

SUBFIG TEN

(d) Subfigure Ten.

Figure 3: Four figures with auto fitting in a minipage.

Under figure/text

(a) First caption.

Under figure/text

(b) Longer second caption.

Under figure/text

(c) Third caption.

Under figure/text

(d) Longer fourth caption. longer fourth caption. longer fourth caption.

Figure 4: Four figures testing caption fitting.